An attacker is trying to target Gandi's Simple Hosting services with a distributed denial of service. Our teams are mitigating the attack. Service speed may be affected. 

 

Update 19:25 CEST We filtered a web hos who was slow to respond to our requests to stop the attackers. The rest of the world can use Simple Hosting normally again. 


We are encountering some technical problems with some Simple Hosting instances. Our team is working on fixing the issue.

Please do not try to restart your instances. We will provide more information on this incident here.

 

15:20 CEST : some virtual machine disks were impacted

15:40 CEST : the problem is fixed, the situation will slowly be returning to normal for the affected virtual machines


We are encountering some problems with an IaaS storage unit. Please do not try to restart your machines. We will provide more information on this below.

 

11:55 : the service is now operating in degraded mode. We will need to change a piece of equipment.


We will be performing a scheduled maintenance on the Gandi Blog database from midnight to 01:00, Paris time (22:00-23:00 GMT) during the night of 23-24 April 2013.

The Blogs will be unreachable during the maintenance.

Sorry for the inconvenience this may cause.

EDIT : 01:08 CEST : maintenance ended


An issue is ongoing on DNS resolvers on the hosting platform.

 

The technical team is analyzing and fixing the issue currently.

Do not reboot your VPS.

 

Update 17:30 : two system administrators are going on site, the rest of the team is triying to restart the DNS resolvers remotely and spawn new DNS resolvers in the meantime.

 

Update 17.35 : one DNS resolver is now working again. You can use :

- 217.70.184.226

- 217.70.184.228

- 217.70.184.230

 

Update 17.45 : the two mostly used DNS resolvers are working again.


We suffer from a DDoS attack on the IaaS Hosting platform on the Paris/FR datacenter.

 

The technical team is currently mitigating the attach and trying to solve the encountered issue.

4 physical nodes have been impacted by the attack.

 

 

UPDATE : End of the issue around 2PM. The technical team did restart the physical nodes and the concerned VMs properly. The attack has been mitigated and is analyzed more precisely now.

 

UPDATE : a IaaS/PaaS storage unit will be rebooted in the next minutes, for safety reaons because of a few disks which are blocked. The reboot operation will take 20 minutes approximately. there is no need to reboot or stop/start your VPS or SH instance, the I/O will be delayed and will catch up once the filer is restarted. We do the necessary operations on our side. Sorry for the inconvenience this issue may cause to you.

 

UPDATE : the reboot operation took less than 10 minutes, the I/O recovered within 5 minutes, the operations are still halted but will be restarted in the next minutes once all is clear.

 

UPDATE : hosting operations are now restarted, the situation is back to normal.



We will be performing major network upgrades to our backbone network in Paris during the night of 23-24 March 2013.  

 

The purpose of this work is to replace our current (aging) backbone network routers with newer generation equipment allowing for increased capacity, performance and resilience.  During this time, all Gandi services in Paris will experience brief periods of network unavailability and/or instability as the migration proceeds.

 

This activity will not impact the individual Gandi services themselves, but rather only the network connectivity between the internet and those services.  As a result, there will be no need for hosting customers to restart their IaaS or PaaS servers/instances.  Services in our Baltimore datacentre will not be directly affected by this maintenance, but some periods of routing instability and reconvergence may be experienced, though these will be far less prevalent than for the Paris services.

 

This migration will commence at 01:30 CET (00:30 GMT) on Sunday 24 March 2013, and we estimate between 10 and 15 minutes of major network outage during the initial migration, followed by a number of brief periods of network instability during the remainder of the migration.  The entire migration activity should last approximately 90 minutes, but no more than two hours.

Operations will be ended at 05:00 AM CET (04:00 AM GMT).

 

We apologise in advance for any inconvenience this may cause.

EDIT 03:08 AM CET : maintenance ended


We will be carrying out maintenance our online documentation at http://wiki.gandi.net today between 14:00 and 15:00 Paris time. Consequently, the interface will not be available during this period.


We are being targeted by DDoS attacks today (starting 13 February around 15h30 CET).

Our infrastructure has experienced two noticeable slowdowns since 15h30, but our teams are handling the attacks as they arrive.

Updates will be provided here.

EDIT 14 feb. 11:36 CET : situation is now stable.


Page   1 2 35 6 711 12 13
Taille du bandeau d'actualités